# Majorization via positivity of Jack and Macdonald polynomial differences

### v1

@@ 3.2. Reductions of the conjectures; proof for two variables
To prove that $\lambda\mathrel{\succcurlyeq}\mu$ implies Eq. 3.1, it suffices to assume that $\lambda$ and $\mu$ are adjacent in the majorization order. In that case, $\lambda=R_{ij}(\mu)$ for some $i<j$, where $R_{ij}(\mu)=(\dots,\mu_{i}+1,\dots,\mu_{j}-1,\dots)$ is the raising operator, see [18, pp. 8].
@@
The first claim follows by transitivity. The second claim is [18, I. (1.16)]. ∎
@@
By the reduction lemmas above, in order to prove the weaker statement $\lambda\mathrel{\succcurlyeq}\mu\implies$ Eq. 3.2, we may assume that
  • either $\lambda=(\nu,k,0)$ and $\mu=(\nu,k-1,1)$, where $\nu\in\mathcal{P}_{n-2}$ such that $\nu_{n-2}\geqslant k\geqslant 2$;
  • or $\lambda=(\nu,2,1^{m},0)$ and $\mu=(\nu,1^{m+2})$, where $1\leqslant m\leqslant n-2$, and $\nu\in\mathcal{P}_{n-m-2}$ such that $\nu_{n-m-2}\geqslant 2$.
The following toy example studies the simplest case of the first kind of these pairs $(\lambda,\mu)$.
Example 3.10 (Toy Example for CGS).
Let $n=2$, $d\geqslant 2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$. Suppressing the argument $\tau$ from $P_{\lambda}(\cdot),P_{\mu}(\cdot)$ henceforth, we have:
$\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}=\sum_{i=0}^{d}\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}x_{1}^{d-i}x_{2}^{i}$ (3.10)
and
$\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}$ $=\frac{x_{1}x_{2}}{1}\frac{P_{\mu-\bm{1}}(x)}{P_{\mu-\bm{1}}(\bm{1})}=\sum_{i=0}^{d-2}\binom{d-2}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i-2}}}{(2\tau)_{{d-2}}}x_{1}^{d-i-1}x_{2}^{i+1}.$ (3.11)
Setting $x=(1,s)$, $s>0$, we have
$f(s)$ $\coloneqq\frac{P_{\lambda}(1,s)}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(1,s)}{P_{\mu}(\bm{1})}$
Define for $0\leqslant i\leqslant d-2$,
$g_{i}\coloneqq\binom{d-2}{i}\frac{(\tau+d-1)}{\tau}\frac{(\tau)_{{i+1}}(\tau)_{{d-i-1}}}{(2\tau)_{{d}}}\in\mathbb{F}_{>0},$
and let $g_{-1}=g_{-2}=g_{d-1}=g_{d}=0$. Then one can show that
$f_{i}=g_{i-2}-2g_{i-1}+g_{i},\text{\quad or equivalently, \quad}\sum_{j=i+1}^{d}(j-i-1)f_{j}=g_{i}.$
This yields Eq. 3.1, because we have for any $s>0$:
$f(s)=(s-1)^{2}g(s)\coloneqq(s-1)^{2}\sum_{i=0}^{d-2}g_{i}s^{i}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}}.$
Combining the results and the example above, Theorem 3.4 follows. ∎
3.3. Okounkov–Olshanski’s formula
As promised above, in this subsection we show Section 3.2. Set $\mathbb{R}^{n}_{\geqslant}\coloneqq\{\,x\in\mathbb{R}^{n}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}x_{1}\geqslant\cdots\geqslant x_{n}\,\}$. For $x\in\mathbb{R}^{n}_{\geqslant}$ and $y\in\mathbb{R}^{n-1}_{\geqslant}$, write $y\prec x$ (interlacing) if
@@ 3.3. Okounkov–Olshanski’s formula
Let $\lambda\in\mathcal{P}_{n-1}$ be a partition of length at most n1n-1, and $(\lambda,0)\in\mathcal{P}_{n}$. Then [21, pp. 77] (where $\theta$ is our $\tau$) shows that for $x\in\mathbb{R}^{n}_{\geqslant}$,
$P_{(\lambda,0)}(x)=\frac{1}{C(\lambda,n;\tau)V(x)^{2\tau-1}}\int_{y\prec x}P_{\lambda}(y)V(y)\Pi(x,y)^{\tau-1}\mathop{}\!\mathrm{d}y,$ (3.13)
@@
We note that Okounkov–Olshanski’s formula has a better form when changed to the following normalization.
@@
$\frac{P_{(\lambda,0)}(x)}{P_{(\lambda,0)}(\bm{1}_{n})}=\int_{y\prec x}\frac{P_{\lambda}(y)}{P_{\lambda}(\bm{1}_{n-1})}K(x,y)\mathop{}\!\mathrm{d}y,\qquad\forall\lambda\in\mathcal{P}_{n-1},\ x\in\mathbb{R}^{n}_{\geqslant}.$ (3.15)
In particular, $\int_{y\prec x}K(x,y)\mathop{}\!\mathrm{d}y=1$.
@@
We are now ready to prove Section 3.2.
Proof of Section 3.2.
Assume that Eq. 3.1 with n1n-1 variables $y=(y_{1},\dots,y_{n-1})$ holds for the pair $(\lambda,\mu)$. We first prove Eq. 3.2 for $x=(x_{1},\dots,x_{n})$ and the pair $((\lambda,0),(\mu,0))$. For $\tau>0$, $x\in\mathbb{R}_{\geqslant 0}^{n}$, since the kernel $K(x,y)>0$, we have
$\frac{P_{(\lambda,0)}(x)}{P_{(\lambda,0)}(\bm{1}_{n})}-\frac{P_{(\mu,0)}(x)}{P_{(\mu,0)}(\bm{1}_{n})}=\int_{y\prec x}\left(\frac{P_{\lambda}(y)}{P_{\lambda}(\bm{1}_{n-1})}-\frac{P_{\mu}(y)}{P_{\mu}(\bm{1}_{n-1})}\right)K(x,y)\mathop{}\!\mathrm{d}y>0.$
Since this is symmetric in $x$, Eq. 3.2 holds for all $x\in[0,\infty]^{n}$. When $\tau=0$, Jack polynomials reduce to monomial symmetric polynomials and Eq. 3.2 is simply Muirhead’s inequality. ∎
@@ 3.4. Supporting evidence: Kadell’s integral
We end this section by providing supporting evidence for the assertion $(4)\implies(2)$ in Conjecture 1. We begin by recalling Kadell’s integral in [18, VI. Ex. 10.7] and [8] (with a different normalization). Namely, let $\lambda$ be a partition of length at most $n$, as usual, and let $r,s>0$ (or more generally, $r,s\in\mathbb{C}$ with $\Re r,\Re s>0$). Then
@@
This supports $(4)\implies(2)$ in Conjecture 1, since for each fixed $\tau=\tau_{0}\in(0,\infty)$, if the difference $\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}$ is positive over $0,1]^{n}$, then so is its integral against Kadell’s density
$\frac{1}{I_{(0^{n})}}|V(x)|^{2\tau}\prod_{i=1}^{n}x_{i}^{r-1}(1-x_{i})^{s-1}.$
@@
Note that $\tilde{I}_{\mu}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$. By Section 3.2, we may assume there exist $1\leqslant i<j\leqslant n$ such that $\lambda=R_{ij}(\mu)$, i.e., $\lambda_{i}=\mu_{i}+1$, $\lambda_{j}=\mu_{j}-1$ and $\lambda_{k}=\mu_{k}$ for $k\neq i,j$. Since $\lambda_{i}-\lambda_{j}-1\geqslant 0$,
@@ 4. Jack polynomial differences and the Muirhead semiring
In this section we prove Conjecture 1 (i.e., that $\lambda\mathrel{\succcurlyeq}\mu\implies$ Eq. 3.1) in some more cases. In fact, we prove a stronger version of the conjecture, in which positivity is replaced by sums of Muirhead differences (which are positive by Muirhead’s inequality).
@@ 4.1. Jack differences in the Muirhead semiring
To remind the reader: consider the outstanding theorem from the introduction (Theorem 1.4) and the ensuing discussion. We upgrade the final statement in that subsection to the following.
Conjecture 3.
Fix $\tau_{0}\in[0,\infty]$. If $\lambda\mathrel{\succcurlyeq}\mu$, then the Jack difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$ (and conversely):
Note that when $\tau_{0}=0$, Conjecture 3 is trivial as $P_{\lambda}(x;0)=m_{\lambda}(x)$. For $\tau_{0}=1$, $P_{\lambda}(x;1)=s_{\lambda}(x)$, and it is (informally) conjectured in [5, Section 7] that the difference in Eq. 4.1 is in the Muirhead cone $\mathcal{M}_{C}(\mathbb{Q}_{\geqslant 0})$. However, we now show that this is false.
Remark 4.1.
Conjecture 3 is false if we use the Muirhead cone instead of the semiring. For example, consider $n=3$, $\lambda=(5,2,0)$, and $\mu=(5,1,1)$. We have
$P(x)$ $\coloneqq 3(2\tau+1)(3\tau+1)(3\tau+2)(3\tau+4)\Bigg(\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\Bigg)$
= $=\;2(\tau+1)(\tau+2)^{2}(2\tau+3)M_{(5,2,0)}-(\tau+1)(\tau+2)(2\tau^{2}+13\tau+12)M_{(5,1,1)}$
$\quad+6\tau(\tau+1)(\tau+2)(2\tau+3)M_{(4,3,0)}-2\tau(\tau+1)(2\tau^{2}+17\tau+17)M_{(4,2,1)}$
$\quad-12\tau(\tau+1)(\tau+2)M_{(3,3,1)}-\tau(\tau+1)(10\tau^{2}+\tau-20)M_{(3,2,2)}.$
If $10\tau^{2}+\tau-20<0$, that is, $0<\tau<\frac{3\sqrt{89}-1}{20}\approx 1.365$, the coefficient of $M_{(3,2,2)}$ is positive. Since $(3,2,2)$ is the minimum (in the majorization order) partition that appears in $P(x)$, $P(x)$ cannot lie in the Muirhead cone. In particular, this happens for the Schur case $\tau=1$. However, this remains consistent with Conjecture 3 because
$P(x)$ $=(2\tau^{4}+19\tau^{3}+43\tau^{2}+54\tau+24)(M_{(5,2,0)}-M_{(5,1,1)})$
@@
which is in the Muirhead semiring $\mathcal{M}_{S}(\mathbb{R}_{\geqslant 0})$ but not in the cone $\mathcal{M}_{C}(\mathbb{R}_{\geqslant 0})$.
@@
Here are three connections of Conjecture 3 to Muirhead’s inequality (Theorem 1.3). First, Conjecture 3, together with the inequality implies Conjecture 1. Second, the converse to this conjecture holds because of Muirhead’s inequality – see Theorem 1.3(2) – and the argument in the proof of Theorem 3.1. Thus Conjecture 3 provides yet another conjectural characterization of majorization. Third, the proof of Conjecture 3 for $\mu=(1^{|\lambda|})$ uses Muirhead’s inequality; see below.
Remark 4.3.
One of the conjectures above ($(4)\implies(1)$ in Conjecture 1) was formulated in all Lie types by McSwiggen and Novak in an interesting paper [19]; see their Conjecture 4.7 and the discussion on pp. 3997. The authors also showed the “easier half” in [19, Proposition 4.8] using asymptotics, as well as the other half for $\dim V=1$ in their Proposition 4.10. While we only work in type $A$, here are four ways in which we go beyond [19]: (a) We show the easier half (see Theorems 3.1 and 3.2) using alternate techniques, which prove the result not just for Jack polynomials but also for power sums (as was suspected in light of Theorem 1.2). These are not Jack specializations, so not covered in [19]. (b) Our method also characterizes weak majorization, again not addressed in [19]. (c) In fact we show Conjecture 3, which is stronger than Conjecture 1 (via the preceding remark) and which is not in [19]. Finally, (d) our proof of the “harder half” holds (for two variables) not just for normalized Jack differences, but also (Theorem 5.6) for Macdonald polynomials, which were “hoped for” in [19].
@@
Conjectures 1, 2 and 3 hold for all $\lambda$ when $\mu=(1^{|\lambda|})$. Conjecture 3 moreover holds when $\lambda$ and $\mu$ are of length at most 2.
Proof of Theorem 4.4.
If $\lambda\mathrel{\succcurlyeq}\mu=(1^{|\lambda|})$, we prove more strongly that the difference lies in the Muirhead cone $\mathcal{M}_{C}(\mathbb{R}_{\geqslant 0})$. Note that $\frac{P_{\lambda}}{P_{\lambda}(\bm{1})}$ is an $\mathbb{F}_{\geqslant 0}$-linear combination of $M_{\nu}$ (see the arguments for [18, VI. (10.15)] for the expansion of monic Jack into monomials; see also [12] for the expansion of integral Jack into augmented monomials), say $\sum_{\nu}u_{\lambda\nu}M_{\nu}$. Then $\sum_{\nu}u_{\lambda\nu}=1$ by our choice of normalization. Also, we have $\frac{P_{\mu}}{P_{\mu}(\bm{1})}=M_{\mu}$. Hence,
$\frac{P_{\lambda}}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}}{P_{\mu}(\bm{1})}=\sum_{\nu}u_{\lambda\nu}(M_{\nu}-M_{\mu})$ (4.2)
lies in the Muirhead cone. This shows Conjecture 3, and hence Conjectures 1 and 2 for $\mu=(1^{|\lambda|})$.
We next come to the case of $\lambda,\mu\in\mathcal{P}_{2}$. Here we continue the calculations in Toy Example 3.2. Expanding the difference of the normalized Jack polynomials in Muirhead differences and suppressing the argument $\tau$ in them as usual, we obtain:
@@
Fix integer partitions $\lambda,\mu\in\mathcal{P}_{2}$ with $\lambda\succcurlyeq\mu$. For $x=(x_{1},x_{2})$, we have
@@
Notice that Section 4.1 does not hold for higher $n$ – not even for $n=3$, even at $\tau=1$, i.e. for Schur polynomials. (This is in parallel to Conjecture 3 also not holding for the Muirhead cone even with $n=3$ and $\tau=1$, as verified in Section 4.1.) For example, working with $(3)\succcurlyeq(1,1,1)$ yields
@@
In fact for two variables, not only does the conjecture imply this assertion, but the converse is also true. This is because – we claim that – every normalized Muirhead-difference $M_{\lambda}(x_{1},x_{2})-M_{\mu}(x_{1},x_{2})$ (for $\lambda\mathrel{\succcurlyeq}\mu$), when divided by $V(x)^{2}=(x_{1}-x_{2})^{2}$, is a Schur-positive polynomial with half-integer coefficients. This claim reduces by transitivity to majorization-adjacent partitions $\lambda=(a+b+1,a),\mu=(a+b,a+1)$ (for $a\geqslant 0,b\geqslant 1$), for which it holds by an explicit computation:

### v2

@@ 3.2.2. Raising operator
To prove Conjecture 1 ($\lambda\mathrel{\succcurlyeq}\mu\Rightarrow$ Eq. 3.1), it suffices to assume that $\lambda$ and $\mu$ are adjacent in the majorization order. In that case, $\lambda=R_{ij}(\mu)$ for some $i<j$, where $R_{ij}(\mu)=(\dots,\mu_{i}+1,\dots,\mu_{j}-1,\dots)$ is the raising operator, see [17, pp. 8].
@@
The first claim follows by transitivity. The second claim is [17, I. (1.16)]. ∎
@@
3.2.3. Appending 0
We first recall the integration formula due to Okounkov–Olshanski [20].
Set $\mathbb{R}^{n}_{\geqslant}\coloneqq\{\,x\in\mathbb{R}^{n}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}x_{1}\geqslant\cdots\geqslant x_{n}\,\}$. For $x\in\mathbb{R}^{n}_{\geqslant}$ and $y\in\mathbb{R}^{n-1}_{\geqslant}$, write $y\prec_{\mathrm{int}}x$ (interlacing) if
@@ 3.2.3. Appending 0
Let $\lambda\in\mathcal{P}_{n-1}$ be a partition of length at most n1n-1, and $(\lambda,0)\in\mathcal{P}_{n}$. Then [20, pp. 77] (where $\theta$ is our $\tau$) shows that for $x\in\mathbb{R}^{n}_{\geqslant}$,
$P_{(\lambda,0)}(x)=\frac{1}{C(\lambda,n;\tau)V(x)^{2\tau-1}}\int_{y\prec_{\mathrm{int}}x}P_{\lambda}(y)V(y)\Pi(x,y)^{\tau-1}\mathop{}\!\mathrm{d}y,$ (3.14)
@@
Okounkov–Olshanski’s formula has a better form when changed to the unital normalization.
@@
$\frac{P_{(\lambda,0)}(x)}{P_{(\lambda,0)}(\bm{1}_{n})}=\int_{y\prec_{\mathrm{int}}x}\frac{P_{\lambda}(y)}{P_{\lambda}(\bm{1}_{n-1})}K(x,y)\mathop{}\!\mathrm{d}y,\qquad\forall\lambda\in\mathcal{P}_{n-1},\ x\in\mathbb{R}^{n}_{\geqslant}.$ (3.16)
In particular, $\int_{y\prec_{\mathrm{int}}x}K(x,y)\mathop{}\!\mathrm{d}y=1$.
@@
Lemma 3.13 (Appending 0).
If Eq. 3.2 holds for the pair $(\lambda,\mu)$ in n1n-1 variables, then Eq. 3.2 holds for the pair $((\lambda,0),(\mu,0))$ in $n$ variables as well.
Proof.
Assume that Eq. 3.1 holds for the pair $(\lambda,\mu)$ in n1n-1 variables $y=(y_{1},\dots,y_{n-1})$. When $\tau=0$, Jack polynomials reduce to monomial symmetric polynomials and Eq. 3.2 is simply Muirhead’s inequality. For every $\tau>0$, $x\in\mathbb{R}_{\geqslant 0}^{n}$, since the kernel $K(x,y)>0$, we have
$\frac{P_{(\lambda,0)}(x)}{P_{(\lambda,0)}(\bm{1}_{n})}-\frac{P_{(\mu,0)}(x)}{P_{(\mu,0)}(\bm{1}_{n})}=\int_{y\prec_{\mathrm{int}}x}\left(\frac{P_{\lambda}(y)}{P_{\lambda}(\bm{1}_{n-1})}-\frac{P_{\mu}(y)}{P_{\mu}(\bm{1}_{n-1})}\right)K(x,y)\mathop{}\!\mathrm{d}y>0.$
Since the LHS is symmetric in $x$, Eq. 3.2 holds for all $x\in[0,\infty]^{n}$. ∎
3.3. An example and the proof
Example 3.14.
Consider $n=2$, $d\geqslant 2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$. Suppressing the parameter $\tau$, we have
$\frac{P_{\lambda}(x_{1},x_{2})}{P_{\lambda}(1,1)}=\sum_{i=0}^{d}\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}x_{1}^{d-i}x_{2}^{i}$ (3.17)
and
$\frac{P_{\mu}(x_{1},x_{2})}{P_{\mu}(1,1)}$ $=\frac{x_{1}x_{2}}{1}\frac{P_{(d-2,0)}(x_{1},x_{2})}{P_{(d-2,0)}(1,1)}=\sum_{i=0}^{d-2}\binom{d-2}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i-2}}}{(2\tau)_{{d-2}}}x_{1}^{d-i-1}x_{2}^{i+1}.$ (3.18)
Setting $x=(1,s)$, $s>0$, we have
$f(s)$ $\coloneqq\frac{P_{\lambda}(1,s)}{P_{\lambda}(1,1)}-\frac{P_{\mu}(1,s)}{P_{\mu}(1,1)}$
Define $g(s)\coloneqq\sum_{i=0}^{d-2}g_{i}s^{i}$, where
$g_{i}\coloneqq\binom{d-2}{i}\frac{(\tau+d-1)}{\tau}\frac{(\tau)_{{i+1}}(\tau)_{{d-i-1}}}{(2\tau)_{{d}}}\in\mathbb{F}_{\geqslant 0},\quad 0\leqslant i\leqslant d-2.$
Let $g_{-1}=g_{-2}=g_{d-1}=g_{d}=0$. Then one can show that
$f_{i}=g_{i-2}-2g_{i-1}+g_{i},$
hence
$f(s)=(s-1)^{2}g(s)\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad s>0.$
Now, Eq. 3.1 follows since
$\frac{P_{\lambda}(x_{1},x_{2})}{P_{\lambda}(1,1)}-\frac{P_{\mu}(x_{1},x_{2})}{P_{\mu}(1,1)}=x_{2}^{d}\cdot f(x_{1}/x_{2})\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad(x_{1},x_{2})\in(0,\infty)^{n}.$
Proof of Theorem 3.2.
In order to prove Conjectures 1 and 2 in the case of two variables, it suffices to prove $\lambda\mathrel{\succcurlyeq}\mu\Rightarrow$ Eq. 3.1 by Theorem 3.4. By Section 3.2.2, it suffices to show this implication for the majorization-adjacent pair: $\lambda=(d+l,l)\mathrel{\succcurlyeq}\mu=(d+l-1,l+1)$, $l\geqslant 0$, $d\geqslant 2$. By Section 3.2.1, it can be further reduced to $\lambda=(d,0)\mathrel{\succcurlyeq}\mu=(d-1,1)$, and this case was proved in Section 3.3.
The next assertion – in which $\mu={\bf 1}_{|\lambda|}$ – is proved in Theorems 4.2 and 4.4 and the remark in between them.
We next show the final assertion. By Section 3.2.3, the weaker implication $(4)\Rightarrow(2)$ in Conjecture 1 holds for partitions $\lambda$ and $\mu$ of length at most 2 and any number of variables. We are now done by Theorem 3.4. ∎
@@ 3.4. Supporting evidence: Kadell’s integral
We end this section by providing supporting evidence for the assertion $(4)\Rightarrow(2)$ in Conjecture 1. We begin by recalling Kadell’s integral in [17, VI. Ex. 10.7] and [7] (with a different normalization). Namely, let $\lambda$ be a partition of length at most $n$, as usual, and let $r,s>0$ (or more generally, $r,s\in\mathbb{C}$ with $\Re r,\Re s>0$). Then
@@
Note that $\tilde{I}_{\mu}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ for $r,s>0$. By Section 3.2.2, we may assume there exist $1\leqslant i<j\leqslant n$ such that $\lambda=R_{ij}(\mu)$, i.e., $\lambda_{i}=\mu_{i}+1$, $\lambda_{j}=\mu_{j}-1$ and $\lambda_{k}=\mu_{k}$ for $k\neq i,j$. Since $\lambda_{i}-\lambda_{j}-1\geqslant 0$,
@@
This supports $(4)\Rightarrow(2)$ in Conjecture 1, since for each fixed $\tau_{0}\in(0,\infty)$, if the difference $\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}$ is positive over $0,1]^{n}$, then so is its integral against Kadell’s density
$\frac{1}{I_{(0^{n})}}|V(x)|^{2\tau_{0}}\prod_{i=1}^{n}x_{i}^{r-1}(1-x_{i})^{s-1}>0.$
@@ 4. Jack polynomial differences and the Muirhead semiring
In this section we prove Conjecture 1 (i.e., that $\lambda\mathrel{\succcurlyeq}\mu\Rightarrow$ Eq. 3.1) in some more cases. In fact, we prove a stronger version of the conjecture, in which evaluation positivity is replaced by sums and products of Muirhead differences (which are positive by Muirhead’s inequality).
@@ 4.1. Jack differences in the Muirhead semiring
Recall the Muirhead cone and semiring, defined above Theorem 1.4.
In [4, Section 7], the authors wondered if normalized Schur differences $\frac{s_{\lambda}(x)}{s_{\lambda}(\bm{1})}-\frac{s_{\mu}(x)}{s_{\mu}(\bm{1})}$ lie in the Muirhead cone, for $\lambda\mathrel{\succcurlyeq}\mu$. However, this is not the case as the example below shows. It seems that the result would hold if one replaces the Muirhead cone by the semiring—which is strictly larger, as the example also confirms.
Example 4.1.
Consider $n=3$, $\lambda=(5,2,0)$, and $\mu=(5,1,1)$. We have
$70\cdot\Bigg(\frac{s_{\lambda}(x)}{s_{\lambda}(\bm{1})}-\frac{s_{\mu}(x)}{s_{\mu}(\bm{1})}\Bigg)=10M_{(5,2,0)}-9M_{(5,1,1)}+10M_{(4,3,0)}-8M_{(4,2,1)}-4M_{(3,3,1)}+M_{(3,2,2)}.$
Note that the coefficient of $M_{(3,2,2)}$ is positive and $(3,2,2)$ is the minimum (in the majorization order) partition that appears in the expansion of $\frac{s_{\lambda}(x)}{s_{\lambda}(\bm{1})}-\frac{s_{\mu}(x)}{s_{\mu}(\bm{1})}$, and hence the difference cannot lie in the Muirhead cone.
However, this difference indeed lies in the Muirhead semiring, in fact for Jack polynomials with arbitrary parameter $\tau=\tau_{0}$. Even more strongly, this holds for the cone $\mathcal{C}=\mathbb{F}_{\geqslant 0}$:
$\mathrel{\phantom{=}}3(2\tau+1)(3\tau+1)(3\tau+2)(3\tau+4)\Bigg(\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\Bigg)$
$=(2\tau^{4}+19\tau^{3}+43\tau^{2}+54\tau+24)(M_{(5,2,0)}-M_{(5,1,1)})$
@@
which is in the Muirhead semiring $\mathcal{M}_{S}(\mathbb{F}_{\geqslant 0})$.
Motivated by this calculation, we make the following conjecture.
Conjecture 3.
The following are equivalent for partitions $\lambda$ and $\mu$:
  1. (1)
    The Jack difference lies in the Muirhead semiring over $\mathbb{F}_{\geqslant 0}$:
    $\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\in\mathcal{M}_{S}(\mathbb{F}_{\geqslant 0}).$ (4.1)
  2. (2)
    For every $\tau_{0}\in[0,\infty]$, the Jack difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$:
  3. (3)
    For some $\tau_{0}\in[0,\infty]$, the Jack difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$:
  4. (4)
    $\lambda$ majorizes $\mu$.
Theorem 4.2.
In Conjecture 3, clearly $(1)\Rightarrow(2)\Rightarrow(3)$, and (3) implies Conjecture 1(3), which gives $\lambda\mathrel{\succcurlyeq}\mu$. So Conjecture 3 reduces to showing $(4)\Rightarrow(1)$. ∎
@@
By Muirhead’s inequality, Conjecture 3(1) implies Conjecture 1(1). Thus Conjecture 3 implies Conjecture 1.
@@
Conjectures 1, 2 and 3 hold for all $\lambda$ when $\mu=(1^{|\lambda|})$. Conjecture 3 moreover holds in two variables.
Proof.
If $\lambda\mathrel{\succcurlyeq}\mu=(1^{|\lambda|})$, we prove more strongly that the difference lies in the Muirhead cone $\mathcal{M}_{C}(\mathbb{F}_{\geqslant 0})$. Write
$\frac{P_{\lambda}}{P_{\lambda}(\bm{1})}=\sum_{\nu\preceq\lambda}u_{\lambda\nu}M_{\nu},\quad u_{\lambda\nu}\in\mathbb{F}_{\geqslant 0}.$
Then $\sum_{\nu}u_{\lambda\nu}=1$ by our choice of normalization. Also, we have $\frac{P_{\mu}}{P_{\mu}(\bm{1})}=M_{\mu}$. Hence,
$\frac{P_{\lambda}}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}}{P_{\mu}(\bm{1})}=\sum_{\nu}u_{\lambda\nu}(M_{\nu}-M_{\mu})\in\mathcal{M}_{C}(\mathbb{F}_{\geqslant 0}).$ (4.4)
This shows Conjecture 3, and hence Conjectures 1 and 2 for $\mu=(1^{|\lambda|})$.
We next come to the case of two variables. Here we continue the calculations in Section 3.3. Expanding the difference of the normalized Jack polynomials in Muirhead differences and suppressing the parameter $\tau$ as usual, we obtain:
@@
Let $x=(x_{1},x_{2})$ and fix $\lambda\mathrel{\succcurlyeq}\mu$ partitions with at most two parts. We have
@@
Notice that Section 4.1 does not hold for higher nn—not even for $n=3$, even at $\tau=1$, i.e. for Schur polynomials. (This is in parallel to Conjecture 3 also not holding for the Muirhead cone even with $n=3$ and $\tau=1$, as verified in Section 4.1.) For example, working with $(3)\mathrel{\succcurlyeq}(1,1,1)$ yields
@@
In fact for two variables, not only does the conjecture imply this assertion, but the converse is also true. This is because—we claim that—every normalized Muirhead-difference $M_{\lambda}(x_{1},x_{2})-M_{\mu}(x_{1},x_{2})$ (for $\lambda\mathrel{\succcurlyeq}\mu$), when divided by $V(x)^{2}=(x_{1}-x_{2})^{2}$, is a Schur-positive polynomial with half-integer coefficients. This claim reduces by transitivity to majorization-adjacent partitions $\lambda=(a+b+1,a),\mu=(a+b,a+1)$ (for $a\geqslant 0,b\geqslant 1$), for which it holds by an explicit computation: